# #1#1# 1 (1)

### v2

@@ Abstract
The pursuit of a generalizable stereo matching model, capable of performing across varying resolutions and disparity ranges without dataset-specific fine-tuning, has revealed a fundamental trade-off. Iterative local search methods achieve high scores on constrained benchmarks, but their core mechanism inherently limits the global consistency required for true generalization. On the other hand, global matching architectures, while theoretically more robust, have been historically rendered infeasible by prohibitive computational and memory costs. We resolve this dilemma with $S^{2}M^{2}$: a global matching architecture that achieves both state-of-the-art accuracy and high efficiency without relying on cost volume filtering or deep refinement stacks. Our design integrates a multi-resolution transformer for robust long-range correspondence, trained with a novel loss function that concentrates probability on feasible matches. This approach enables a more robust joint estimation of disparity, occlusion, and confidence. $S^{2}M^{2}$ establishes a new state of the art on the Middlebury v3 and ETH3D benchmarks, significantly outperforming prior methods across most metrics while reconstructing high-quality details with competitive efficiency.
@@ 1 Introduction
Recent advancements in deep learning have significantly improved stereo matching accuracy, enabling high-quality depth estimation in well-constrained conditions [4, 31, 1, 32, 20, 18, 42, 37, 23]. However, ensuring reliable performance across diverse environments and image resolutions remains challenging. A robust stereo matching model should generalize across different conditions (e.g., indoor/outdoor, low/high texture scenes) without requiring dataset-specific tuning. This pursuit for generalization often involves increasing model capacity for improved accuracy, but such gains do not always scale proportionally with size, demanding carefully designed architectures and training strategies. Furthermore, larger models inherently bring higher computational costs, necessitating a balance between efficiency and strong generalization, especially at high resolutions .
The landscape of deep learning-based stereo matching is primarily defined by a trade-off between two dominant architectural paradigms: global matching and iterative refinement. Transformer-based global matching architectures [19, 45, 38, 39, 23] are inherently well-suited for modeling long-range correspondences to handle occlusions and repetitive patterns. However, they have been historically underexplored because they pose significant efficiency challenges and are difficult to scale, both in processing high-resolution images and in effectively increasing model capacity. This scalability bottleneck led to the dominance of cost volume-based iterative refinement methods [20, 49, 18, 42], which demonstrated strong performance on established benchmarks. This paradigm, however, faces two fundamental limitations. First, processing the initial cost volume requires computationally expensive filtering, such as 3D convolutions, which severely hampers scalability. Second, the subsequent iterative refinement, being local in nature, struggles to resolve complex and discontinuous disparity structures, often failing to escape local minima.
@@
Figure 2: Comprehensive evaluation on two representative benchmarks, ETH3D [29] and Middlebury v3 [28]. Accuracy is measured using the Bad-2.0 metric for Middlebury and Bad-0.5 for ETH3D. Each method is plotted with its corresponding parameter size.
@@ 2 Related Works
Figure 3: Our proposed architecture, S2M², comprises four main stages: (1) feature extraction, where left and right input images are processed by FPN and MRT to yield multi-resolution feature maps; (2) global matching, which computes all-pairs correlation at 1/4 resolution and estimates disparity, occlusion, and confidence using optimal transport; (3) refinement, iteratively updating these estimates; and (4) upsampling, enhancing depth edges to restore full resolution. The feature extraction stage employs a Multi-Resolution Transformer (MRT) for hierarchical feature extraction, incorporating an Adaptive Gated Fusion Layer (AGFL) for efficient multi-scale feature alignment.
@@
In parallel, cost volume-based iterative refinement has emerged as a dominant pathway, demonstrating strong practical performance. This approach was pioneered by RAFT-Stereo [20], which introduced recurrent updates on a learned correlation volume. Subsequent works have advanced this framework through several key strategies. For instance, CREStereo [18] introduced a more robust, hierarchical recurrent refinement to better handle fine details and real-world image distortions. Another line of work, such as IGEV [42], incorporated geometric guidance to improve disparity propagation. More recently, efforts have centered on refining feature representations, with Selective-IGEV [37] selectively integrating frequency information and MoCha-Stereo [5, 6] employing motif-based attention for robustness. Despite these advancements, the performance of these methods hinges on a high-quality correlation volume produced by computationally expensive filtering, and the inherently local nature of this process makes them less effective in globally ambiguous regions with repetitive textures or large discontinuities, reinforcing the need for true global correspondence mechanisms.
@@
A recent trend aims to improve generalization by leveraging large-size, pre-trained models. These methods typically build upon the iterative refinement framework but replace the standard feature encoder with a powerful backbone pre-trained on other large-scale tasks. This is often achieved by incorporating a general-purpose vision encoder or even a complete monocular depth estimation model as a feature extractor [50, 40, 16, 3]. By transferring rich prior knowledge, these approaches demonstrate impressive generalization capabilities. However, this strategy introduces a significant trade-off: their reliance on massive, often fixed-weight components results in models that are too large and computationally expensive for many practical applications. This highlights the ongoing challenge of achieving both strong generalization and efficiency, a balance our scalable architecture aims to achieve.
@@ 3.1.1 Multi-Resolution Transformer (MRT)
The MRT is designed to overcome the efficiency limitations of conventional Transformers. Instead of processing a single high-resolution feature map, it operates on the entire feature pyramid in parallel. Its efficiency stems from a hybrid attention strategy: to align with epipolar geometry, attention is restricted to the horizontal axis (1D) for higher resolutions, while full 2D self-attention is applied only at the coarsest level ($\mathbf{F}_{32}$) to capture global context. This design strikes a critical balance between performance and computational cost.
The internal structure of each MRT block, illustrated in Figure 3, is also carefully designed. Each block is composed of a self-attention and a cross-attention module, with each respective attention layer being followed by an FFN to produce highly discriminative features. To ensure stable training of a deep stack of MRT blocks, each block is also equipped with cross-scale skip connections.
@@ 3.1.2 Adaptive Gated Fusion Layer (AGFL)
To build a high-capacity feature extractor, we stack multiple MRT blocks. A key challenge in such a deep, multi-resolution architecture is ensuring effective information exchange and coherent learning across all scales. Prior architectures like U-Net can suffer from gradient bottlenecks at the highest resolution, hindering the training of coarser-level features [23]. While our parallel MRT architecture avoids this specific bottleneck, it still requires a mechanism for robust cross-scale communication.
@@ 4.1 Probabilistic Mode Concentration (PMC) Loss
Global matching is performed on 1/4 downsampled feature maps, resulting in at most 16 disparity candidates at each pixel in this lower resolution space. Let $K$ (where $K\leq 16$) denote the number of these candidates. For each candidate $c_{ij}^{(k)}$, we define its neighborhood as
@@
Iters GFLOPs Disparity Metrics Auxiliary Metrics
@@ 6 Conclusion
This paper presents S²M², a multi-resolution Transformer framework that overcomes the long-standing challenges of global matching, positioning it as a practical, state-of-the-art methodology. The proposed framework is highly scalable, robustly handling challenging high resolutions and wide disparity ranges, while its performance improves consistently with model size. It produces reliable depth estimates by jointly modeling disparity, confidence, and occlusion. Our extensive experiments not only establish a new state-of-the-art on real-world benchmarks but also demonstrate the framework’s robustness in demanding scenarios, thus opening new avenues for developing highly accurate and generalizable stereo systems.

### v3

@@ Abstract
The pursuit of a generalizable stereo matching model, capable of performing well across varying resolutions and disparity ranges without dataset-specific fine-tuning, has revealed a fundamental trade-off. Iterative local search methods achieve high scores on constrained benchmarks, but their core mechanism inherently limits the global consistency required for true generalization. However, global matching architectures, while theoretically more robust, have historically been rendered infeasible by prohibitive computational and memory costs. We resolve this dilemma with S²M²: a global matching architecture that achieves state-of-the-art accuracy and high efficiency without relying on cost volume filtering or deep refinement stacks. Our design integrates a multi-resolution transformer for robust long-range correspondence, trained with a novel loss function that concentrates probability on feasible matches. This approach enables a more robust joint estimation of disparity, occlusion, and confidence. S²M² establishes a new state of the art on Middlebury v3 and ETH3D benchmarks, significantly outperforming prior methods in most metrics while reconstructing high-quality details with competitive efficiency.
@@ 1 Introduction
Recent advances in deep learning have significantly improved stereo matching accuracy, enabling high-quality depth estimation in well-constrained conditions [4, 31, 1, 32, 20, 18, 42, 37, 23]. However, ensuring reliable performance across diverse environments and image resolutions remains a challenge. A robust stereo matching model should generalize across different conditions (e.g., indoor/outdoor, low/high texture scenes) without requiring dataset-specific tuning. This pursuit of generalization often involves increasing model capacity for improved accuracy, but such gains do not always scale proportionally with size, requiring carefully designed architectures and training strategies. Furthermore, larger models inherently bring higher computational costs, necessitating a balance between efficiency and strong generalization, especially at high resolutions.
The landscape of deep learning-based stereo matching is primarily defined by a trade-off between two dominant architectural paradigms: global matching and iterative refinement. Transformer-based global matching architectures [19, 45, 38, 39, 23] are particularly well-suited for modeling long-range correspondences to handle occlusions and repetitive patterns. However, they have historically been underexplored because they pose significant efficiency challenges and are difficult to scale, both in processing high-resolution images and in effectively increasing model capacity. This scalability bottleneck led to the dominance of cost volume based iterative refinement methods [20, 49, 18, 42], which demonstrated strong performance on established benchmarks. This paradigm, however, faces two fundamental limitations. First, processing the initial cost volume requires computationally expensive filtering, such as 3D convolutions, which severely hampers scalability. Second, the subsequent iterative refinement, being local in nature, struggles to resolve complex and discontinuous disparity structures, often failing to escape local minima.
@@
Figure 2: Comprehensive evaluation on two representative benchmarks, ETH3D [29] and Middlebury v3 [28]. Accuracy is measured using the Bad-2.0 metric for Middlebury and Bad-0.5 for ETH3D. Each method is plotted with its corresponding model size (in millions of parameters).
@@ 2 Related Works
Figure 3: Our proposed architecture, S2M², comprises four main stages: (1) feature extraction, where left and right input images are processed by FPN and MRT to yield multi-resolution feature maps; (2) global matching, which computes all-pairs correlation at 1/4 resolution and estimates disparity, occlusion, and confidence using optimal transport; (3) refinement, iteratively updating these estimates; and (4) upsampling, enhancing depth boundaries to restore full resolution. The feature extraction stage employs a Multi-Resolution Transformer (MRT) for hierarchical feature extraction, incorporating an Adaptive Gated Fusion Layer (AGFL) for efficient multi-scale feature alignment.
@@
In parallel, cost volume based iterative refinement has emerged as a dominant pathway, demonstrating strong practical performance. This approach was pioneered by RAFT-Stereo [20], which introduced recurrent updates on a learned correlation volume. Subsequent works have advanced this framework through several key strategies. For instance, CREStereo [18] introduced a more robust, hierarchical recurrent refinement to better handle fine details and real-world image distortions. Another line of work, such as IGEV [42], incorporated geometric guidance to improve disparity propagation. More recently, efforts have centered on refining feature representations, with Selective-IGEV [37] selectively integrating frequency information and MoCha-Stereo [5, 6] employing motif-based attention for robustness. Despite these advancements, the performance of these methods hinges on a high-quality correlation volume produced by computationally expensive filtering, and the inherently local nature of this process makes them less effective in globally ambiguous regions with repetitive textures or large discontinuities, reinforcing the need for true global correspondence mechanisms.
@@
A recent trend aims to improve generalization by leveraging large-size, pre-trained models. These methods typically build upon the iterative refinement framework but replace the standard feature encoder with a powerful backbone pre-trained on other large-scale tasks. This is often achieved by incorporating a general purpose vision encoder or even a complete monocular depth estimation model as a feature extractor [50, 40, 16, 3]. By transferring rich prior knowledge, these approaches demonstrate impressive generalization capabilities. However, this strategy introduces a significant trade-off: their reliance on massive, often fixed-weight components results in models that are too large and computationally expensive for many practical applications. This highlights the ongoing challenge of achieving both strong generalization and efficiency, a balance our scalable architecture aims to achieve.
@@ 3.1.1 Multi-Resolution Transformer (MRT)
The MRT is designed to overcome the efficiency limitations of conventional Transformers. Instead of processing a single high-resolution feature map, it operates on the entire feature pyramid in parallel. Its efficiency stems from a hybrid attention strategy: to align with epipolar geometry, attention is restricted to the horizontal axis (1D) at higher resolutions, while full 2D self-attention is applied only at the coarsest level ($\mathbf{F}_{32}$) to capture global context. This design strikes a critical balance between performance and computational cost.
The internal structure of each MRT block, illustrated in Figure 3, is also carefully designed. Each block is composed of a self-attention module and a cross-attention module, with each respective attention layer being followed by an FFN to produce highly discriminative features. To ensure stable training of a deep stack of MRT blocks, each block is also equipped with cross-scale skip connections.
@@ 3.1.2 Adaptive Gated Fusion Layer (AGFL)
We build a high-capacity feature extractor by stacking multiple MRT blocks. A key challenge in such a deep, multi-resolution architecture is ensuring effective information exchange and coherent learning across all scales. Prior architectures like U-Net can suffer from gradient bottlenecks at the highest resolution, hindering the training of coarser-level features [23]. While our parallel MRT architecture avoids this specific bottleneck, it still requires a mechanism for robust cross-scale communication.
@@ 4.1 Probabilistic Mode Concentration (PMC) Loss
Global matching is performed on 1/4 downsampled feature maps, resulting in at most 16 disparity candidates at each pixel in this lower resolution space. Let $K$ (with $K\leq 16$) denote the number of these candidates. For each candidate $c_{ij}^{(k)}$, we define its neighborhood as
@@
Iterations GFLOPs Disparity Metrics Auxiliary Metrics
@@ 6 Conclusion
This paper presents S²M², a multi-resolution Transformer framework that overcomes the long-standing challenges of global matching, positioning it as a practical, state-of-the-art methodology. The proposed framework is highly scalable, robustly handling challenging high resolutions and wide disparity ranges, while its performance improves consistently with model size. It produces reliable depth estimates by jointly modeling disparity, confidence, and occlusion. Our extensive experiments not only establish a new state-of-the-art on real-world benchmarks but also demonstrate the framework’s robustness in demanding scenarios, thereby opening new avenues for developing highly accurate and generalizable stereo systems.